# KernJC: Automated Vulnerable Environment Generation for Linux Kernel Vulnerabilities

### v2

@@ 1. Introduction
Vulnerabilities often occur in specific subsystems or as a result of particular module features, which may not be active by default. Nevertheless, this important detail is frequently omitted in both CVE databases and external reports, complicating the process of activating kernel vulnerabilities. What’s worse, the reliability of config files from successful reproduction attempts is also questionable, as evidenced by sources like (bsauce, 2022a, b) and our observations of vulnerability disclosure (ANGUELKOV, [n. d.]; Krysiuk, [n. d.]) in § 5.5.1. Similar to the situation in (Chen et al., 2021), it is not practical to enable all kernel configs as well, as some configs are mutually exclusive, and others may affect the execution of target functionality. Although there is an allyesconfig option available in the Linux kernel build mechanism enabling as many configs as possible, our manual experiments and prior work (Hao et al., 2023; You et al., 2017) confirm that the kernel built with allyesconfig is not bootable.
@@ 2.2. Challenges
Specifically, the three red lines from the vulnerable function xt_compat_target_from_user in Figure 2 (a), which are deleted by the patch (Westphal, 2022) in Figure 2 (b), have already disappeared in the v5.11.22 source code (Harald Welte, [n. d.]).
@@
In terms of CVE-2021-22555, our analysis of a segment of its patch (Westphal, 2022), as illustrated in Figure 2 (b), identifies crucial configs in the Makefiles along the path leading to the inclusion of net/netfilter/x_tables.c. These include CONFIG_NETFILTER and CONFIG_NETFILTER_XTABLES in Figure 2 (c), as well as two other configs, CONFIG_NET and CONFIG_INET, upon which the former two configs depend, as specified in the Kconfig files in Figure 2 (d). Furthermore, within the vulnerable file itself in Figure 2 (a), a conditional compilation directive (#ifdef) is observed, controlling the inclusion of the vulnerable function based on CONFIG_COMPAT. By conducting a thorough manual analysis of all files affected by the patch in this manner, we can compile a comprehensive list of configs necessary to ensure the presence of the vulnerability in the target kernel version. The detailed analysis results for CVE-2021-22555, including the essential configs, are presented in Figure 2 (e).
Unfortunately, even with the heuristic analysis above, the PoC for CVE-2021-22555 (Google, 2022) remains non-functional. The issue stems from the PoC’s specification of NFQUEUE as the target in Netfilter, depicted in a code snippet from the PoC in Figure 2 (f). This additionally requires the activation of CONFIG_NETFILTER_XT_TARGET_NFQUEUE to support the specified target. Ultimately, enabling this config allows the PoC to trigger the vulnerability successfully.
@@
Vulnerability descriptions and reports often lack detailed information regarding the necessary configs that would ensure both the presence and user space accessibility of the vulnerability. Analysts have to download the source code of a vulnerable version, then compile the kernel using default configs (make defconfig) and execute the PoC program to ascertain if it either crashes the system or triggers a Kernel Address SANitizer (KASAN) report.
This approach is generally effective for vulnerabilities located in the kernel’s core functionalities, which are usually enabled by default configs. However, for numerous vulnerabilities tied to non-default functionalities or intricate subsystems, such as CVE-2021-22555, relying solely on default configs proves inadequate for their activation. A manual analytical approach is typically needed to identify additional required configs. Alternatively, a heuristic analysis of the vulnerable code is often used alongside an examination of the Kbuild (kernel.org, 2024a) and Kconfig (kernel.org, 2014) mechanisms. Despite its monolithic nature, the kernel source code is hierarchically organized and developed modularly. The Kconfig and Kbuild systems work in tandem to tailor the kernel’s functionalities and facilitate its build process. Statistically, the latest Linux kernel encompasses over 15,000 config items (CONFIG_*), the majority of which are instrumental in deciding whether to incorporate specific source code files and activate particular features. Typically, kernel vulnerabilities manifest at the code block level. Thus, pinpointing the essential configs for introducing a vulnerability entails selecting the relevant configs from these 15,000 items, ensuring the inclusion of files and code blocks associated with the vulnerability. Given the extensive number of items and their intricate interrelations, manually undertaking this task is notably labor-intensive. What’s worse, some configs needed by the reproduction are determined by analyzing the PoC programs, which are not always readily available, particularly for vulnerabilities that have only recently been disclosed.
@@ 3.1. Overview
Environment Provisioning. After KernJC discerns the kernel configs essential for the vulnerability, it integrates these configs with the foundational ones (defconfig). Following this integration, KernJC proceeds to compile the kernel’s source code of the identified vulnerable version, to build the kernel image. This constructed image, in conjunction with a root filesystem (rootfs), is then utilized to provision a virtual machine as the final reproduction environment atop hypervisors.
@@ 3.3. Vulnerability Version Identification
This specific patch incorporates both a line deletion and a line addition, framed within a 3-line context (represented by the grey rectangles around the target line in Figure 6). The first patching action substitutes the original if statement with a new one, thereby remedying the vulnerability. Consequently, a subsequent patch attempt is unsuccessful, as the target line for deletion is no longer present following context alignment. This indicates that the source code has already undergone patching.
Essentially, the patch commit(s) content is derived from the diff output for affected files in kernel source code of old & new versions (kernel.org, 2024c). The finest granularity of patches is at the line level, with differences being delineated on a per-line basis, e.g., line additions or line deletions. The patch for a Linux kernel vulnerability could be composed of one or more commits; each commit may contain diff results for one or more files; each file’s diff result contains either line additions, line deletions, or both. However, only the record of line modifications is inadequate for patching. Location information should also be provided to locate the accurate position to be patched. Due to the activity of Linux kernel development, locating via line numbers is prone to be erroneous when other commits occur in the same file. Linux kernel community adopts a N-line context way (Juhl, 2024) to provide location information, i.e., attaching N lines before and after the modified lines as context into the patch, which can tolerate the aforementioned line shift phenomenon to some extent. Although sometimes it would still fail (e.g., when both the N-line context and the affected line(s) are replicated in the same file), this approach is easy to implement and efficient.
@@ 3.4. Vulnerability Config Identification
The specification of this directed graph is presented in Table 1. There are three types of config definition in the Kconfig mechanism: config, menuconfig, and inner config within blocks (e.g., choice block), which are represented by vertexes in our graph.
Additionally, menu is a special item in Kconfig, which also has relations with other configs, but it is not a real config.
@@
For example, as shown in § 2.2, the menu "Core Netfilter Configurations" depends on the following three configs: CONFIG_NET, CONFIG_INET, and CONFIG_NETFILTER.
As a result, there will be three depend edges from the virtual vertex "Core Netfilter Configurations" to the three config vertexes in the graph. Similarly, CONFIG_INET has an opaque_depend edge to CONFIG_NET, as the former config is defined in the if block of the latter one. The select and imply edges represent the relations in Kconfig that when config A is enabled, it will select or imply config B to be enabled as well, if there is a select CONFIG_B statement within the definition of CONFIG_A.
@@
Thirdly, KernJC leverages the direct configs and the Kconfig graph to get the hidden configs, which includes three types of configs (vertexes in the graph): (1) configs that are recursively reachable from any direct config (), (2) configs holding one-hop select relation to any direct config (), and (3) configs holding one-hop depend relation to any direct config (). As edges in the Kconfig graph are directed, the reachability of serves as a strong attribute to identify the vulnerability configs, which is consequently handled recursively. Nonetheless, and hold inverse relations to direct configs, which are only used to fulfill the config requirement for functionalities in complex subsystems, so KernJC conducts one-hop discovery to avoid introducing too many weakly related configs. Lastly, the set , composed of both direct configs and hidden configs, serves as the final result for the vulnerability config identification process. It is noticeable that virtual vertexes are not considered as one hop, as they do not represent real configs, and will be filtered out from . In summary, the identified configs can be categorized into six categories, as shown in Figure 8.
@@
For CVE-2021-22555, the heuristic analysis result in Figure 2 (e) is equal to the combination of , and derived by the graph-based approach. Especially, the last pivotal config to activate this vulnerability, CONFIG_NETFILTER_XT_TARGET_NFQUEUE, is successfully captured in . Therefore, this approach can provide adequate configs to support the reproduction of CVE-2021-22555, illustrated by the result graph shown in Figure 9. Furthermore, § 5.3 will analyze the contributions of configs in different sets (, , , , and ) to the reproduction of representative kernel vulnerabilities.
@@
The approach to identifying necessary kernel configs for a specific vulnerability is derived from two observations: (1) to include a vulnerability, the vulnerable code must be activated by conditional compilation like #ifdef (code-level configs), and the directory path towards the vulnerable file(s) must be included in Makefile(s) along the path (path-level configs); (2) the code-level and path-level configs have multiple necessary relations with other configs in the hierarchical Kconfig mechanism. Respectively, we regard configs in (1) as direct configs, in that they are intuitive and can be found out by analyzing the patch and source code, and the configs in (2) as hidden configs, because they hide in the complex structure of Kconfig and have to be identified with deep analysis. As a result, our initial idea to identify necessary configs is to first figure out all code-level () and path-level () configs, build a directed graph for all kernel configs and their relations from Kconfig files in the vulnerable source code, and use these configs as starting points to discover all reachable configs as hidden configs ().
@@
However, only configs above do not necessarily ensure the activation of one vulnerability. There are exceptions in the reproduction experiences. For example, to trigger CVE-2017-18344 (NVD, 2017), the PoC must access a pseudo-file, /proc/pid/timers, which is only available when CONFIG_CHECKPOINT_RESTORE is set (Kerrisk, [n. d.]). For this case, the CVE description in NVD rarely provides the needed config items. Besides, for vulnerabilities located in complicated subsystems (e.g., Netfilter), special configs within the subsystems are needed to ensure the normal functionality in user space, illustrated by the CVE-2021-22555 example in § 2.2. Consequently, , , and are proposed to improve the identification result. Considering the complexity of Linux Kconfig mechanism, we only derive and from one-hop relations, instead of two or more hops, to avoid involving too many useless configs. This strategy proves to be effective and adequate in § 5.3.
@@ 4. Implementation
Patch Processing. Patches play a crucial role in identifying both the presence of vulnerabilities and the necessary configs. KernJC integrates crawlers for both the Ubuntu Security Tracker (Ubuntu, 2024) and Red Hat Bugzilla (Hat, 2024), aimed at collecting patch commit IDs related to kernel vulnerabilities. When required, KernJC utilizes these IDs to fetch the raw patch content from the official kernel repository (kernel.org, 2024b). The system then parses this data to pinpoint the modified files and specific line changes. These files are subsequently retrieved for applying patches and executing a direct config search. For identifying configs, KernJC inspects each vulnerable function’s end where modifications occur, tracing back to the file’s beginning to detect any conditional compilation directives (#ifdef CONFIG_*).
Kconfig Graph Construction. The creation of a Kconfig graph is intricate due to the complex relations within the Linux kernel’s Kconfig system. For example, a source "net/packet/Kconfig" line within an if NET condition implies that all configs in the net/packet/Kconfig file are dependent on CONFIG_NET. KernJC tackles this by applying a recursive strategy to form graph edges that represent these relations. Technologically, the tool utilizes NetworkX (NetworkX, 2014) to construct the graph and support traversal analysis.
Rootfs and Virtual Machine Setup. KernJC employs an adapted version of the create-image script from the syzkaller project (Google, 2024b) for generating the base image for rootfs. In constructing each vulnerable environment, KernJC uses QEMU’s (kashyapc.fedorapeople.org, [n. d.]) copy-on-write overlay feature to quickly establish rootfs for the targeted virtual machine (VM). Finally, the tool engages QEMU to launch the virtual machine, facilitating the vulnerability reproduction process.
User-friendly Interface. The final phase in vulnerability reproduction is the testing of the developed PoC within the execution environment. This traditionally involves a cumbersome process where the analyst must transfer the PoC into the environment, execute it, and then monitor its behavior. Success on the initial attempt is rare, leading to a repetitive cycle of development and debugging of the PoC until it functions as intended. To streamline this often arduous task, KernJC’s command-line interface has been thoughtfully designed to resemble Docker’s user-friendly interface. It includes commands like cp for copying files between the host and the VM, exec for executing commands in the VM, and attach for interacting with the VM, all aimed at simplifying the delivery and execution of the PoC within this VM.
@@ 5.2. Performance in Reproduction
For each vulnerability in the 66-CVE dataset, we run KernJC to build the vulnerable environments and then compile and execute the associated PoC in the target environments for reproduction. For each vulnerability with the source code downloaded, we run KernJC twice to build two environments with different config identification approaches: (1) for the first one, KernJC builds the environment with the identified configs using the aforementioned graph-based approach, and (2) for the second one, KernJC builds it with the default configs (make defconfig) as baseline. The PoC will be executed in both of these two environments. After the PoC is executed, we observe whether a KASAN report is generated in the kernel log or the special OS state described in the PoC file is achieved, to determine the result of reproduction.
We conduct extra reproduction experiments for each vulnerability with false positive version claims detected to ensure the correctness of the version identification process. Specifically, we run KernJC to build a reproduction environment with a detected false positive version and KernJC-identified configs (a combination of identified non-default configs and the default configs from make defconfig) for each case and execute the PoC within it.

### v3

@@ 1. Introduction
Vulnerabilities often occur in specific subsystems or as a result of particular module features, which may not be active by default. Nevertheless, this important detail is frequently omitted in both CVE databases and external reports, complicating the process of activating kernel vulnerabilities. What’s worse, the reliability of config files from successful reproduction attempts is also questionable, as evidenced by sources like (bsauce, 2022a, b) and our observations of vulnerability disclosure (ANGUELKOV, [n. d.]; Krysiuk, [n. d.]) in § 5.5.1. Similar to the situation in (Chen et al., 2021), it is not practical to enable all kernel configs as well, as some configs are mutually exclusive, and others may affect the execution of target functionality. Although there is an allyesconfig
option available in the Linux kernel build mechanism enabling as many configs as possible, our manual experiments and prior work (Hao et al., 2023; You et al., 2017) confirm that the kernel built with allyesconfig
is not bootable.
@@ 2.2. Challenges
Specifically, the three red lines from the vulnerable function xt_compat_target_from_user
in Figure 2 (a), which are deleted by the patch (Westphal, 2022) in Figure 2 (b), have already disappeared in the v5.11.22 source code (Harald Welte, [n. d.]).
@@
In terms of CVE-2021-22555, our analysis of a segment of its patch (Westphal, 2022), as illustrated in Figure 2 (b), identifies crucial configs in the Makefiles along the path leading to the inclusion of net/netfilter/x_tables.c
. These include CONFIG_NETFILTER
and CONFIG_NETFILTER_XTABLES
in Figure 2 (c), as well as two other configs, CONFIG_NET
and CONFIG_INET
, upon which the former two configs depend, as specified in the Kconfig files in Figure 2 (d). Furthermore, within the vulnerable file itself in Figure 2 (a), a conditional compilation directive (#ifdef
) is observed, controlling the inclusion of the vulnerable function based on CONFIG_COMPAT
. By conducting a thorough manual analysis of all files affected by the patch in this manner, we can compile a comprehensive list of configs necessary to ensure the presence of the vulnerability in the target kernel version. The detailed analysis results for CVE-2021-22555, including the essential configs, are presented in Figure 2 (e).
Unfortunately, even with the heuristic analysis above, the PoC for CVE-2021-22555 (Google, 2022) remains non-functional. The issue stems from the PoC’s specification of NFQUEUE
as the target in Netfilter, depicted in a code snippet from the PoC in Figure 2 (f). This additionally requires the activation of CONFIG_NETFILTER_XT_TARGET_NFQUEUE
to support the specified target. Ultimately, enabling this config allows the PoC to trigger the vulnerability successfully.
@@
Vulnerability descriptions and reports often lack detailed information regarding the necessary configs that would ensure both the presence and user space accessibility of the vulnerability. Analysts have to download the source code of a vulnerable version, then compile the kernel using default configs (make defconfig
) and execute the PoC program to ascertain if it either crashes the system or triggers a Kernel Address SANitizer (KASAN) report.
This approach is generally effective for vulnerabilities located in the kernel’s core functionalities, which are usually enabled by default configs. However, for numerous vulnerabilities tied to non-default functionalities or intricate subsystems, such as CVE-2021-22555, relying solely on default configs proves inadequate for their activation. A manual analytical approach is typically needed to identify additional required configs. Alternatively, a heuristic analysis of the vulnerable code is often used alongside an examination of the Kbuild (kernel.org, 2024a) and Kconfig (kernel.org, 2014) mechanisms. Despite its monolithic nature, the kernel source code is hierarchically organized and developed modularly. The Kconfig and Kbuild systems work in tandem to tailor the kernel’s functionalities and facilitate its build process. Statistically, the latest Linux kernel encompasses over 15,000 config items (CONFIG_*
), the majority of which are instrumental in deciding whether to incorporate specific source code files and activate particular features. Typically, kernel vulnerabilities manifest at the code block level. Thus, pinpointing the essential configs for introducing a vulnerability entails selecting the relevant configs from these 15,000 items, ensuring the inclusion of files and code blocks associated with the vulnerability. Given the extensive number of items and their intricate interrelations, manually undertaking this task is notably labor-intensive. What’s worse, some configs needed by the reproduction are determined by analyzing the PoC programs, which are not always readily available, particularly for vulnerabilities that have only recently been disclosed.
@@ 3.1. Overview
Environment Provisioning. After KernJC discerns the kernel configs essential for the vulnerability, it integrates these configs with the foundational ones (defconfig
). Following this integration, KernJC proceeds to compile the kernel’s source code of the identified vulnerable version, to build the kernel image. This constructed image, in conjunction with a root filesystem (rootfs), is then utilized to provision a virtual machine as the final reproduction environment atop hypervisors.
@@ 3.3. Vulnerability Version Identification
This specific patch incorporates both a line deletion and a line addition, framed within a 3-line context (represented by the grey rectangles around the target line in Figure 6). The first patching action substitutes the original if
statement with a new one, thereby remedying the vulnerability. Consequently, a subsequent patch attempt is unsuccessful, as the target line for deletion is no longer present following context alignment. This indicates that the source code has already undergone patching.
Essentially, the patch commit(s) content is derived from the diff
output for affected files in kernel source code of old & new versions (kernel.org, 2024c). The finest granularity of patches is at the line level, with differences being delineated on a per-line basis, e.g., line additions or line deletions. The patch for a Linux kernel vulnerability could be composed of one or more commits; each commit may contain diff
results for one or more files; each file’s diff
result contains either line additions, line deletions, or both. However, only the record of line modifications is inadequate for patching. Location information should also be provided to locate the accurate position to be patched. Due to the activity of Linux kernel development, locating via line numbers is prone to be erroneous when other commits occur in the same file. Linux kernel community adopts a N-line context way (Juhl, 2024) to provide location information, i.e., attaching N lines before and after the modified lines as context into the patch, which can tolerate the aforementioned line shift phenomenon to some extent. Although sometimes it would still fail (e.g., when both the N-line context and the affected line(s) are replicated in the same file), this approach is easy to implement and efficient.
@@ 3.4. Vulnerability Config Identification
The specification of this directed graph is presented in Table 1. There are three types of config definition in the Kconfig mechanism: config
, menuconfig
, and inner config
within blocks (e.g., choice
block), which are represented by vertexes in our graph.
Additionally, menu
is a special item in Kconfig, which also has relations with other configs, but it is not a real config.
@@
For example, as shown in § 2.2, the menu "Core Netfilter Configurations"
depends on the following three configs: CONFIG_NET
, CONFIG_INET
, and CONFIG_NETFILTER
As a result, there will be three depend
edges from the virtual vertex "Core Netfilter Configurations"
to the three config vertexes in the graph. Similarly, CONFIG_INET
has an opaque_depend
edge to CONFIG_NET
, as the former config is defined in the if
block of the latter one. The select
and imply
edges represent the relations in Kconfig that when config A is enabled, it will select
or imply
config B to be enabled as well, if there is a select CONFIG_B
statement within the definition of CONFIG_A
@@
Thirdly, KernJC leverages the direct configs and the Kconfig graph to get the hidden configs, which includes three types of configs (vertexes in the graph): (1) configs that are recursively reachable from any direct config (), (2) configs holding one-hop select
relation to any direct config (), and (3) configs holding one-hop depend
@@
relation to any direct config (). As edges in the Kconfig graph are directed, the reachability of serves as a strong attribute to identify the vulnerability configs, which is consequently handled recursively. Nonetheless, and hold inverse relations to direct configs, which are only used to fulfill the config requirement for functionalities in complex subsystems, so KernJC conducts one-hop discovery to avoid introducing too many weakly related configs. Lastly, the set , composed of both direct configs and hidden configs, serves as the final result for the vulnerability config identification process. It is noticeable that virtual vertexes are not considered as one hop, as they do not represent real configs, and will be filtered out from . In summary, the identified configs can be categorized into six categories, as shown in Figure 8.
@@
For CVE-2021-22555, the heuristic analysis result in Figure 2 (e) is equal to the combination of , and derived by the graph-based approach. Especially, the last pivotal config to activate this vulnerability, CONFIG_NETFILTER_XT_TARGET_NFQUEUE
, is successfully captured in . Therefore, this approach can provide adequate configs to support the reproduction of CVE-2021-22555, illustrated by the result graph shown in Figure 9. Furthermore, § 5.3 will analyze the contributions of configs in different sets (, , , , and ) to the reproduction of representative kernel vulnerabilities.
The approach to identifying necessary kernel configs for a specific vulnerability is derived from two observations: (1) to include a vulnerability, the vulnerable code must be activated by conditional compilation like #ifdef
@@
(code-level configs), and the directory path towards the vulnerable file(s) must be included in Makefile(s) along the path (path-level configs); (2) the code-level and path-level configs have multiple necessary relations with other configs in the hierarchical Kconfig mechanism. Respectively, we regard configs in (1) as direct configs, in that they are intuitive and can be found out by analyzing the patch and source code, and the configs in (2) as hidden configs, because they hide in the complex structure of Kconfig and have to be identified with deep analysis. As a result, our initial idea to identify necessary configs is to first figure out all code-level () and path-level () configs, build a directed graph for all kernel configs and their relations from Kconfig files in the vulnerable source code, and use these configs as starting points to discover all reachable configs as hidden configs ().
However, only configs above do not necessarily ensure the activation of one vulnerability. There are exceptions in the reproduction experiences. For example, to trigger CVE-2017-18344 (NVD, 2017), the PoC must access a pseudo-file, /proc/pid/timers
, which is only available when CONFIG_CHECKPOINT_RESTORE
@@
is set (Kerrisk, [n. d.]). For this case, the CVE description in NVD rarely provides the needed config items. Besides, for vulnerabilities located in complicated subsystems (e.g., Netfilter), special configs within the subsystems are needed to ensure the normal functionality in user space, illustrated by the CVE-2021-22555 example in § 2.2. Consequently, , , and are proposed to improve the identification result. Considering the complexity of Linux Kconfig mechanism, we only derive and from one-hop relations, instead of two or more hops, to avoid involving too many useless configs. This strategy proves to be effective and adequate in § 5.3.
@@ 4. Implementation
Patch Processing. Patches play a crucial role in identifying both the presence of vulnerabilities and the necessary configs. KernJC integrates crawlers for both the Ubuntu Security Tracker (Ubuntu, 2024) and Red Hat Bugzilla (Hat, 2024), aimed at collecting patch commit IDs related to kernel vulnerabilities. When required, KernJC utilizes these IDs to fetch the raw patch content from the official kernel repository (kernel.org, 2024b). The system then parses this data to pinpoint the modified files and specific line changes. These files are subsequently retrieved for applying patches and executing a direct config search. For identifying configs, KernJC inspects each vulnerable function’s end where modifications occur, tracing back to the file’s beginning to detect any conditional compilation directives (#ifdef CONFIG_*
Kconfig Graph Construction. The creation of a Kconfig graph is intricate due to the complex relations within the Linux kernel’s Kconfig system. For example, a source "net/packet/Kconfig"
line within an if NET
condition implies that all configs in the net/packet/Kconfig
file are dependent on CONFIG_NET
. KernJC tackles this by applying a recursive strategy to form graph edges that represent these relations. Technologically, the tool utilizes NetworkX (NetworkX, 2014) to construct the graph and support traversal analysis.
Rootfs and Virtual Machine Setup. KernJC employs an adapted version of the create-image
script from the syzkaller project (Google, 2024b) for generating the base image for rootfs. In constructing each vulnerable environment, KernJC uses QEMU’s (kashyapc.fedorapeople.org, [n. d.]) copy-on-write overlay feature to quickly establish rootfs for the targeted virtual machine (VM). Finally, the tool engages QEMU to launch the virtual machine, facilitating the vulnerability reproduction process.
User-friendly Interface. The final phase in vulnerability reproduction is the testing of the developed PoC within the execution environment. This traditionally involves a cumbersome process where the analyst must transfer the PoC into the environment, execute it, and then monitor its behavior. Success on the initial attempt is rare, leading to a repetitive cycle of development and debugging of the PoC until it functions as intended. To streamline this often arduous task, KernJC’s command-line interface has been thoughtfully designed to resemble Docker’s user-friendly interface. It includes commands like cp
for copying files between the host and the VM, exec
for executing commands in the VM, and attach
for interacting with the VM, all aimed at simplifying the delivery and execution of the PoC within this VM.
@@ 5.2. Performance in Reproduction
For each vulnerability in the 66-CVE dataset, we run KernJC to build the vulnerable environments and then compile and execute the associated PoC in the target environments for reproduction. For each vulnerability with the source code downloaded, we run KernJC twice to build two environments with different config identification approaches: (1) for the first one, KernJC builds the environment with the identified configs using the aforementioned graph-based approach, and (2) for the second one, KernJC builds it with the default configs (make defconfig
) as baseline. The PoC will be executed in both of these two environments. After the PoC is executed, we observe whether a KASAN report is generated in the kernel log or the special OS state described in the PoC file is achieved, to determine the result of reproduction.
We conduct extra reproduction experiments for each vulnerability with false positive version claims detected to ensure the correctness of the version identification process. Specifically, we run KernJC to build a reproduction environment with a detected false positive version and KernJC-identified configs (a combination of identified non-default configs and the default configs from make defconfig
) for each case and execute the PoC within it.